OpenStudy (anonymous):

Factor the following: 1. 25a2 + 40x + 16 2. 6y2 + 5y + 1 3. 20z2 + 22z - 12 4. a2 + 3ab - 4b2 PLEASE HELP & EXPLAIN HOW TO DO THEM

OpenStudy (anonymous):

fo number two i would us "AC Method"

OpenStudy (anonymous):

okay, explain what that is

OpenStudy (anonymous):

so here it goes: to use this method, the expression has to be in AX^2 + Bx + C form okaii.

OpenStudy (anonymous):

get it until now?

OpenStudy (anonymous):

i dont want to loose you early?

OpenStudy (anonymous):

yess!! that makes complete since, the a,b and c will be filled in with the number in the question. lol: a is 6, b is 5,c is 1, okay go(:

OpenStudy (anonymous):

ok good :)

OpenStudy (anonymous):

Steps: 1.) multiply AC 2.) Find 2 integers such that the product equals AC and the sum equals B 3.) Rewrite the Original problem by using A,C, and the 2 new integers. 4.) Factor by using Grouping.

OpenStudy (anonymous):

r u still good until now?

OpenStudy (anonymous):

1.) would that just stay 6 cuz 6*1=6 or would it be 12 because u do 6^2*1

OpenStudy (anonymous):

or i mean 6^2*1=36 lol

OpenStudy (anonymous):

no, ur right :0 it would be 6

OpenStudy (anonymous):

okayy! then i get confused on 2.)

OpenStudy (anonymous):

keep going :)

OpenStudy (anonymous):

the 5y+1= 6. is that what u mean?

OpenStudy (anonymous):



OpenStudy (anonymous):

get it:)

OpenStudy (anonymous):

so for step 2, you got the 3 because 3 goes into 6 twice? right? then step 3 u just rewrote the problem with the new equation and step 4 u grouped everything. yes?

OpenStudy (anonymous):

haha im alittle confused on the last problem u drew

OpenStudy (anonymous):

for step 2 i used those numbers (3 and 2) because when multiplied it equals 6, right? but when you add 3 + 2 = 5 hence the A = 6 AND B = 5

OpenStudy (anonymous):

okay, i get it(: just couldnt explain what i was thinking

OpenStudy (anonymous):

i think the GCF of all the things in #1 is one so im guessing it can only be factored to 1(25a2+40x+10) That doesnt really go down that much so i may be wrong

OpenStudy (anonymous):

@JaneDoe100 ur correct that is what i told her earlier but i deleted my post, to not get her confused. i would say that it is prime, dont you think?

OpenStudy (anonymous):

so it would be a=4,3 because 4*-1=-4 but 4+-1=3?

OpenStudy (anonymous):

a=-4,3*

OpenStudy (anonymous):

OHHHHH, yesssss i seee, i see.

OpenStudy (anonymous):

ok lets start over on # 4, cause i read it wrong on ur question lol sorry :(

OpenStudy (anonymous):

a^(2)+3ab-4b^(2) In this problem 4*-1=-4 and 4-1=3, so insert 4 as the right hand term of one factor and -1 as the right-hand term of the other factor. (a+4b)(a-b)

OpenStudy (anonymous):